What is y to the power of 0
Answer by stanbon(75887)

What is y to the power of 0
-----
y^0 = 1 as long as y is not zero.
=======================================
Why is that?
2/2 = 2^(1-1) = 2^0
-----
But 2/2 = 1
-----
Therefore, 2^0 = 1
======================
cheers,
Stan H.
========================